OpenStudy (anonymous):

if kinetic energy is increases by 300% then change in momentum is 100% 200% 300%

OpenStudy (anonymous):

\[ K_i = \frac{1}{2}mv_i^2 = \frac{p_i^2}{2m} \] \[K_f = \frac{p_f^2}{2m} \\ p_f = np_i \\K_f = 3K_i \] \[3K_i =3\frac{p_i^2}{2m} = \frac{(np_i)^2}{2m} \\ 3p_i^2 = n^2 p_i^2 \\n = \sqrt 3 \] The change in the square of the momentum is 300%, but the change in the momentum itself is \[ (100\sqrt3) \% \] :)

OpenStudy (vincent-lyon.fr):

Remember that "increased by 300%" means "multiplied by 4".

OpenStudy (anonymous):

well fiddle. You're right.

OpenStudy (anonymous):

substituting 4 for 3, the last expression is then \[ 4p_i^2 = n^2p_i^2 \] \[ n = 2\]

OpenStudy (anonymous):

so if you double the momentum, you quadruple the kinetic energy!

OpenStudy (anonymous):

HOW Kf = 3 Ki ?? @AllTehMaffs MASTER :)

OpenStudy (anonymous):

I misread the original question - it should be Kf = 4 Ki ^_^

OpenStudy (anonymous):

But from where we get 4 ???

OpenStudy (anonymous):

from the 300% increase. If you have energy K, then 300% increase = 4K so you have an initial K, Ki, and final K, Kf Ki = K Kf = 4K Kf = 4Ki

OpenStudy (anonymous):

How 300 % = 4 ??? SIR :/

OpenStudy (anonymous):

start with K 100% increase = 2K 200% increase = 3K 300% increase = 4K (I got it wrong, too :P)

OpenStudy (anonymous):

Its ok ..:) So NOw what is answer ??

OpenStudy (anonymous):

n = 2 ; pf = 2pi increased by 100% percent The post two above your first comment ^^